OpenStudy (anonymous):

Determine whether the following series converges or diverges using the integral test.

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty \infty}\frac{ 1 }{ k^2+4 }\]

OpenStudy (amistre64):

integrate it then

OpenStudy (amistre64):

let k = 2u dk = 2 du and it should arctan up

OpenStudy (anonymous):

How can I show that? Is it also convergent?